# Fixation and extinction in time-fluctuating spatially structured metapopulations

### v1

@@ IV.0.1 Weak bottlenecks: $\psi(m,K_{-})\gg 1$
The density $p_{\xi}(\nu,n)$ has support $[K_{-},K_{+}]$, and the normalisation constants $\mathcal{Z}_{\pm}$ ensure $\int_{K_{-}}^{K_{+}}p_{\xi}(\nu,n)~{}\text{d}n=1$. The $M/W$ competition characterising the intermediate switching regime dynamics can be described by the coarse-grained representation of Sec. III.0.1 generalised to a time-fluctuating environment following Refs. [16, 17, 22]. Here, we analyse the influence of $\nu$ and $m$ on the $M$ fixation probability, $\Phi_{i}^{\rm G}(\nu,m)$, and uMFT, $\Theta_{i}^{\rm G}(\nu,m)$, in a metapopulation consisting of $i$ mutants demes and $\Omega-i$ $W$-demes spatially arranged as a regular graph ${\rm G}$. To this end, we consider a birth-death process for the number $i=0,\dots,\Omega$ of $M$ demes forming an $M$-cluster. As in Sec. III.0.1, we assume that there is initially a single $M$ deme ($i=1$). The effective rates for increase or decrease in size of the $M$-cluster, $\mathcal{T}^{\pm}_{i}$, depend on the expected number of migrating cells, which in turn depends on the deme size that is now a time-fluctuating quantity driven by Eq. (5). In a time-varying environment, the expected number of migrants from a deme, $mn$, is approximated by $m\mathcal{N}_{\xi}(\nu)$, where the the long-time mean deme size in the environmental state $\xi$ is obtained from the PDMP density according to
$\mathcal{N}_{\xi}(\nu)=\int_{K_{-}}^{K_{+}}np_{\xi}(\nu/s,n)~{}\text{d}n,$ (18)
@@
where, as in Refs. [16, 17, 18, 19, 23], the switching rate has been rescaled, $\nu\to\nu/s$, by the timescale of the deme fixation dynamics, see Appendix B, where there are an average of ${\cal O}(\nu/s)$ switches on the deme fixation timescale [17, 18]. The (marginal) average deme size regardless of $\xi$ is given by $\mathcal{N}(\nu)=\frac{1}{2}\sum_{\xi}\mathcal{N}_{\xi}(\nu)$ and known to be a decreasing function of $\nu$ [16, 17]. As in static environments, see Eq. (10), the transition rates $\mathcal{T}^{\pm}_{i}$ depend on the spatial structure, via $E_{{\rm G}}(i)/q_{{\rm G}}$, and on the probability $\rho_{M/W,\xi}(\nu)$ that an $M/W$ migrant invades a $W/M$ deme in the environment $\xi$. Putting everything together, this yields the effective transition rates
@@
With Eq. (19), by dropping all explicit dependence except on $i$ and $\xi$, we obtain the $M$ fixation probability in the environmental state $\xi$, denoted by $\Phi_{i,\xi}^{{\rm G}}(\nu,m,K_{\pm})$, as the solution of the following $\nu$-dependent first-step analysis equation:
@@
subject to the boundary conditions $\Phi_{0,\xi}^{{\rm G}}=0$ and $\Phi_{\Omega,\xi}^{{\rm G}}=1$. The metapopulation uMFT in the environmental state $\xi$, denoted by $\Theta_{i,\xi}^{{\rm G}}$, satisfies a similar equation:
@@
$\Phi_{i}^{\rm G}(\nu,m)$ $=\frac{1}{2}\sum_{\xi}\Phi_{i,\xi}^{{\rm G}}(\xi,\nu,m),$ (23)
$\Theta_{i}^{\rm G}(\nu,m)$ $=\frac{1}{2}\sum_{\xi}\Theta_{i,\xi}^{{\rm G}}(\xi,\nu,m),$
@@
Eq. (24) are the expressions of the $M$ fixation probability and metapopulation uMFT in the realm of the combined coarse-grained and PDMP description. This approach is valid under slow migration rate ($m\ll 1$) and weak selection strength ($s\ll 1$) for the assumption $n\approx\mathcal{N}_{\xi}(\nu)$ to hold at each invasion, see Appendix A.2. In Fig. 4, the comparison of the predictions of Eq. (24) with the simulation results of the full model on the regular graphs ${\rm G}=\{{\rm clique,cycle,grid}\}$ shows that (24) captures well the dependence of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ on $\nu$, $m$, $s$ and $K_{+}$. In particular, Eq. (24) reproduces on all ${\rm G}$ the non-monotonic $\nu$-dependence of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ (when it exhibits this feature), as well as their behaviour when $\nu\to 0,\infty$ given by Eqs. (14) and (16).
@@
Another distinctive feature of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ is their non-monotonic $\nu$-dependence when the other parameters ($s$, $m$, $K_{\pm}$, $\Omega$) are kept fixed. In particular, Fig. 4 shows that $\Phi^{\rm G}$ may exhibit a sharp peak in the regime of intermediate $\nu\in[10^{-4},10^{-1}]$ that is well captured by Eq. (24). These results are in marked contrast with their counterparts in a single deme, which vary monotonically with $\nu$ [16, 17, 19]. The non-monotonic $\nu$-dependence $\Phi^{\rm G}$ and $\Theta^{\rm G}$ is therefore an inherent effect of spatial migration. Intuitively, this behaviour stems, on the one hand, from more $M$ invasions occurring when the deme size is as close as possible to $n\approx K_{+}$, see Eq. (19). On the other hand, the average deme size is a decreasing function of $\nu$ [16, 17, 18], see Appendix A.2. Therefore, optimising the probability of $M$ fixation requires two considerations: the environment should avoid remaining stuck in the harsh environment for too long, which can happen with a probability close to $1/2$ when $\nu\ll 1$; and the environment should not switch too frequently (i.e. $\nu\gg 1$), as this reduces the effective deme size ($n\approx{\cal K}$), which can be significantly smaller than $K_{+}$. Hence, the best conditions for the fixation of $M$ are for a range of $\nu$ in the intermediate regime. Since the uMFT is longer in the harsh than in the mild environment (where there are fewer migration events, see Fig. 2(b)), a similar reasoning leads to a minimum mean fixation time for $\nu$ in the intermediate regime.
IV.0.2 Strong bottlenecks: $\psi(m,K_{-})<1$
@@
Figure 5: Typical single realisations of $N/\Omega$ (black), $N_{M}/\Omega$ (red), $N_{W}/\Omega$ (blue), and $K(t)$ (grey) vs time for different values of $K_{-}$ and $\nu$. (a,b): Here, $\nu=10^{-4}$ and $K_{-}=8$. In (a), $K=K_{-}$ at $t=0$ and $M$ and then $W$ quickly go extinct. In (b), $K=K_{+}$ at $t=0$ and $M$ fixes the population while $W$ goes extinct. (c,d): Here, $\nu=10^{-2}$ and $K_{-}=8$. In (c), mutants survive the first few bottlenecks but their abundance is low leading to the fixation of $W$ and removal of $M$ after four bottlenecks ($t\gtrsim 1000$). In (d), mutants survive the first bottlenecks and spread in the mild state where they recolonise and invade demes. They are eventually able to fix the population. (e,f): Here, $\nu=10$, and $K_{-}=4$ in (e) and $K_{-}=10$ in (f). $K(t)$ switches very frequently and is not shown for clarity. In (e), the deme size is $n\approx 2K_{-}=8$ and the dynamics is dominated by deme extinction leading to the rapid extinction of the metapopulation. In (e), the deme size is $n\approx 2K_{-}=20$ and there is $M/W$ competition that leads to fixation of $M$ and extinction of $W$ after a typical time $t\sim\theta^{\rm clique}(2K_{-})\gtrsim 10^{4}$ (not shown). Similar results are obtained on other regular graphs ${\rm G}$, see text. Other parameters are $\Omega=10$, $s=0.1$, $m=10^{-4}$, and $K_{+}=200$. In all panels, initially there is a single $M$ deme and $\Omega-1$ demes occupied by $W$.
@@
Figure 6: Near-optimal condition for the idealised treatment. $(\nu,K_{-})$ heatmaps of $\Phi$, $\Theta$, $\Theta_{E}$ and $(1-\Phi)(1-\Theta/\Theta_{E})$ for a clique metapopulation. The migration rate is $m=10^{-3}$ in (a-d) and $m=10^{-4}$ in (e-h). Whitespace in panels (c) and (g) indicate where at least one realisation for those parameters did not reach extinction by $t=10^{5}$, i.e. $\Theta_{E}\sim 10^{5}$ or larger. Grey lines in panels (d) and (h) show the near-optimal conditions for the idealised treatment: $\psi(m,K_{-})<1$ below the top horizontal line, $mK_{+}\theta_{E}>1$ above the bottom horizontal line, and $\nu\theta_{E}>1$ above the curved line, while the vertical line indicates where $\nu<1$ and $\theta_{E}$ from Eq. (13). The near-optimal treatment conditions is the yellowish cloud at the centre of the area enclosed by these lines. Similar results are obtained on other regular graphs ${\rm G}$, see text and Fig. 10. Other parameters $\Omega=16$, $s=0.1$, and $K_{+}=200$. In all panels, initially there is a single $M$ deme and $\Omega-1$ demes occupied by $W$.
@@
When $\psi(m,K_{-})<1$, each deme can undergo strong bottlenecks, see Fig. 3(a) and below. In the harsh environment $\xi=-1$ ($K=K_{-}$), the entire metapopulation experiences extinction, in an observable time $\theta_{E}(K_{-},\Omega)$, denoted $\theta_{E}$ here for conciseness. However, in the mild state $\xi=1$ ($K=K_{+}$), deme extinction can be neglected and each site can be regarded as being occupied by either $W$ or $M$ cells. The dynamics in the harsh state is thus dominated by extinction, whereas the $M/W$ competition characterising the mild state is aptly captured by the two-state coarse-grained description of Section III.0.1. Environmental switching thus couples regimes that are dominated in turn by deme extinction and $M/W$ competition, yielding complex dynamical scenarios whose analysis is difficult. However, we can gain valuable insight by considering first the limits $\nu\ll 1$, $\nu\gg 1$, and then the case of intermediate switching.
@@
When the environment varies very slowly, $\nu\ll 1$, the $K(t)$ remains at its initial value for long periods, that is $K=K_{\pm}$ if $\xi(0)=\pm 1$ each with a probability $1/2$. On the one hand, if initially $\xi=-1$ (harsh environment), $K=K_{-}$ and each deme is prone to extinction after a mean time $\tau_{E}(K_{-})$, which eventually leads to the collapse of the metapopulation after a mMET $\theta_{E}\approx e^{K_{-}}\ln{(\Omega)}/K_{-}$ when $\Omega\gg 1$ and $K_{-}\gg 1$, see Eqs. (6) and (13) and Fig. 5(a). On the other hand, if initially $\xi=1$ (mild condition), $n\approx K_{+}$ and there is $M/W$ competition characterised by the fixation of $M$ with a probability $\phi(K_{+})$ approaching 1 when $K_{+}s\gg 1$, see Eq. (12) and Fig. 5(b). As a result, when $\nu\ll 1$ and $K_{+}s\gg 1$, there are two equally likely outcomes illustrated in Fig. 5: either the extinction of the metapopulation in a mean time $\theta_{E}$ as in Fig. 5(a), or the fixation of $M$ after a mean time $\theta^{\rm G}(K_{+})$ as in Fig. 5(b).
@@
In frequently varying environments, when $\nu\gg 1$, the size of each deme readily settles about the effective carrying capacity Eq. (15), with $n\approx{\cal K}$ after $t\sim 1$ [16, 17]. Since ${\cal K}\approx 2K_{-}$ when $K_{+}\gg K_{-}$, if $\psi(m,2K_{-})<1$ the dynamics is characterised by the extinction of individual demes and then of the entire metapopulation after mean times $\tau_{E}(2K_{-})$ and $\theta_{E}(\Omega,2K_{-})$, see Fig. 5(e). However, if $\psi(m,2K_{-})\gg 1$ and $2K_{-}s\gg 1$, the dynamics is characterised by $M/W$ competition with $M$ most likely to fix the metapopulation after a mean time $\theta^{\rm G}(2K_{-})$, as illustrated by Fig. 5(f).
@@
In the intermediate switching regime, $\nu\lesssim 1$ with $\psi(m,K_{-})<1$, the metapopulation experiences strong bottlenecks, can avoid extinction for extended periods of time, and either strain can prevail. In the harsh environmental state ($K=K_{-}$), the dynamics is always dominated by deme extinction. In the mild state ($K=K_{+}\gg K_{-}$), there is recolonisation of empty demes that rapidly become either $W$ or $M$ demes, followed by invasions and $M/W$ competition. In order to ensure that the collapse of the metapopulation is unlikely to be observed, the mean time spent in either environmental state needs to be shorter than the metapopulation mean extinction time in the harsh environment, i.e. $1/\nu<\theta_{E}$. Moreover, when $\nu\sim 1/\tau_{E}(K_{-})>1/\theta_{E}$, numerous demes go extinct in the harsh environment before switching to the mild state. Hence, when $\nu\lesssim 1$ and $\nu\theta_{E}>1$, the metapopulation is unlikely to go extinct and transiently consists of a mixture of empty demes and $W/M$ demes before either $W$ or $M$ eventually takes over. In this regime, mutants are likely to be removed from the metapopulation when there is a small fraction of them in the harsh environment, see Fig. 5(c). At each strong bottleneck, $M$ demes have a finite probability to go extinct before switching to the mild environment, where surviving mutants can invade $W$ demes and recolonise empty demes. In a scenario illustrated by Fig. 5(c), there are periods of duration $\sim 1/\nu$ during which the number of mutants remains low and prone to extinction when $K=K_{-}$, followed by periods in $K_{+}$ where the number of $M$ demes increases (due to $M/W$ competition). Each bottleneck can thus be regarded as an attempt to remove $M$ demes, whereas each switch $K_{-}\to K_{+}$ can be envisioned as a rescue of mutants. This cycle repeats itself until $M$ demes are entirely removed after enough bottlenecks. The metapopulation thus consists of a fluctuating number of $W$ demes and empty demes. This scenario is the most likely to occur when the initial fraction of $M$ demes is small. Another possible outcome, illustrated by Fig. 5(d), occurs when mutants surviving the harsh conditions invade and are successful in recolonising many demes in the mild environment. Mutants can thus significantly increase the number of $M$ demes, exceeding that of $W$ demes. In this case, bottlenecks can be seen as attempt to remove $W$ and $M$ demes, and the most likely outcome is the removal of $W$ demes. In this scenario, the metapopulation eventually consists of a fluctuating number of empty demes and mutant demes, as illustrated by Fig. 5(d). The results of Fig. 5 have been obtained for cliques, but the same qualitative behaviour is expected for any regular graphs ${\rm G}$, with the spatial structure affecting the the long-term fraction of occupied demes and therefore the probability of removal of each bottleneck. However, phenomena operated by extinction are mostly independent of ${\rm G}$ and $m$, as illustrated by Fig. 10.
@@
A hypothetical idealised treatment: In this intermediate switching regime, the metapopulation is likely to avoid extinction in the harsh environment if $\nu\theta_{E}\gtrsim 1$. Moreover, when $mK_{+}\theta_{E}\gtrsim 1$ enough demes are recolonised in the mild environment to ensure that the metapopulation will not readily go extinct after a bottleneck. Hence, the metapopulation is likely to avoid extinction when $\nu\theta_{E}>1$ and $mK_{+}\theta_{E}>1$ and either $W$ or $M$ can be entirely removed, with respective probabilities $\Phi$ and $1-\Phi$, after a mean time $\Theta$, whereas the mMET in the time-fluctuating environment, $\Theta_{E}$, occurs on a much longer timescale ($\Theta_{E}\gg\Theta$). As an application, we consider a hypothetical idealised treatment to efficiently remove deleterious mutants from an otherwise healthy system by controlling the environmental conditions via the parameters $K_{-}$ and $\nu$. In this context, $M$ cells are interpreted as deleterious mutants that have a selective advantage over healthy $W$ cells composing an organism, here represented by the metapopulation consisting initially of $\Omega-1$ demes of type $W$ and a single $M$ deme. Small numbers of pathogenic cells, as few as 10 for some strains of Salmonella and Shigella [86, 87], can cause illness and so these initial conditions are relevant, even for smaller systems. The idealised treatment consists of finding the set of near-optimal environmental conditions to remove $M$ cells and minimise the risk of extinction of the entire metapopulation. This corresponds to determining the range of $K_{-}$ and $\nu$ for which $\Phi$ and $\Theta/\Theta_{E}$ are minimal. According to the above discussion, the near-optimal conditions for this idealised treatment on a regular graph ${\rm G}$ are

### v2

@@ IV.1 Weak bottlenecks: $\psi(m,K_{-})\gg 1$
The density $p_{\xi}(n;\nu)$ has support $K_{-},K_{+}]$, and the normalisation constants $\mathcal{Z}_{\pm}$ ensure $\int_{K_{-}}^{K_{+}}p_{\xi}(n;\nu)~\text{d}n=1$. The $M/W$ competition characterising the intermediate switching regime dynamics can be described by the coarse-grained representation of Sec. III.1 generalised to a time-fluctuating environment following Refs. [16, 17, 22]. Here, we analyse the influence of $\nu$ and $m$ on the $M$ fixation probability, $\Phi_{i}^{\rm G}(\nu,m)$, and uMFT, $\Theta_{i}^{\rm G}(\nu,m)$, in a metapopulation consisting of $i$ mutants demes and $\Omega-i$ $W$-demes spatially arranged as a regular graph $\rm G$}. To this end, we consider a birth-death process for the number $i=0,\dots,\Omega$ of $M$ demes. As in Sec. III.1, we assume that there is initially a single $M$ deme ($i=1$). The effective rates, denoted by $\mathcal{T}^{\pm}_{i,\xi}$, for the increase or decrease by one of the number $i$ of $M$ demes in the environmental state $\xi$ depend on the expected number of migrating cells, which in turn depends on the deme size that is now a time-fluctuating quantity driven by Eq. (6). In a time-varying environment, the expected number of migrants from a deme, $mn$, is approximated by $m\mathcal{N}_{\xi}(\nu)$, where the the long-time mean deme size in the environmental state $\xi$ is obtained from the PDMP density according to
$\mathcal{N}_{\xi}(\nu)=\int_{K_{-}}^{K_{+}}np_{\xi}(n;\nu/s)~\text{d}n,$ (20)
@@
where, as in Refs. [16, 17, 18, 19, 23], the switching rate has been rescaled, $\nu\to\nu/s$, by the timescale of the deme fixation dynamics (see Appendix A.2) where there are an average of ${\cal O}(\nu/s)$ switches on the deme fixation timescale [17, 18]. The (marginal) average deme size regardless of $\xi$ is given by $\mathcal{N}(\nu)=\frac{1}{2}\sum_{\xi}\mathcal{N}_{\xi}(\nu)$ and known to be a decreasing function of $\nu$ [16, 17]. As in static environments (see Eq. (11)) the transition rates $\mathcal{T}^{\pm}_{i}$ depend on the spatial structure, via $E_{{\rm G}}(i)/q_{{\rm G}}$, and on the probability $\rho_{M/W,\xi}(\nu)$ that an $M/W$ migrant invades a $W/M$ deme in the environment $\xi$. Putting everything together, this yields the effective transition rates
@@
With Eq. (21), by dropping all explicit dependence except on $i$ and $\xi$, we obtain the $M$ fixation probability starting from the environmental state $\xi$ with $i$ mutant demes on a graph $\rm G$, denoted by $\Phi_{i,\xi}^{{\rm G}}(\nu,m)$, as the solution of the following $\nu$-dependent first-step analysis equation:
@@
subject to the boundary conditions $\Phi_{0,\xi}^{{\rm G}}=0$ and $\Phi_{\Omega,\xi}^{{\rm G}}=1$. The metapopulation uMFT starting from the same initial conditions, denoted by $\Theta_{i,\xi}^{{\rm G}}$, satisfies a similar equation:
@@
$\Phi_{i}^{\rm G}(\nu,m)$ $=\frac{1}{2}\sum_{\xi}\Phi_{i,\xi}^{{\rm G}}(\nu,m),$ (25)
$\Theta_{i}^{\rm G}(\nu,m)$ $=\frac{1}{2}\sum_{\xi}\Theta_{i,\xi}^{{\rm G}}(\nu,m),$
@@
Eq. (26) are the expressions of the $M$ fixation probability and metapopulation uMFT in the realm of the combined coarse-grained and PDMP description. This approach is valid under slow migration rate ($m\ll 1$) and weak selection strength ($s\ll 1$) for the assumption $n\approx\mathcal{N}_{\xi}(\nu)$ to hold at each invasion; see Appendix A.2. In Fig. 4, the comparison of the predictions of Eq. (26) with the simulation results of the full model on the regular graphs ${\rm G}=\{{\rm clique,cycle,grid}\}$ shows that Eq. (26) captures well the dependence of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ on $\nu$, $m$, $s$ and $K_{+}$. In particular, Eq. (26) reproduces on all $\rm G$} the non-monotonic $\nu$-dependence of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ (when it exhibits this feature), as well as their behaviour when $\nu\to 0,\infty$ given by Eqs. (16) and (18).
@@
Another distinctive feature of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ is their non-monotonic $\nu$-dependence when the other parameters ($s$, $m$, $K_{\pm}$, $\Omega$) are kept fixed. In particular, Fig. 4 shows that $\Phi^{\rm G}$ may exhibit a sharp peak in the regime of intermediate $\nu\in[10^{-4},10^{-1}]$ that is well captured by Eq. (26). These results are in marked contrast with their counterparts in a single deme, which vary monotonically with $\nu$ [16, 17, 19]. The non-monotonic $\nu$-dependence of $\Phi^{\rm G}$ and $\Theta^{\rm G}$ is therefore an inherent effect of spatial migration. Intuitively, this behaviour stems, on the one hand, from more $M$ invasions occurring when the deme size is as close as possible to $n\approx K_{+}$; see Eq. (21). On the other hand, the average deme size is a decreasing function of $\nu$ [16, 17, 18]; see Appendix A.2. Therefore, optimising the probability of $M$ fixation requires two considerations: the environment should avoid remaining stuck in the harsh environment for too long, which can happen with a probability close to $1/2$ when $\nu\ll 1$; and the environment should not switch too frequently (i.e. $\nu\gg 1$), as this reduces the effective deme size ($n\approx{\cal K}$), which can be significantly smaller than $K_{+}$. Hence, the best conditions for the fixation of $M$ are for a range of $\nu$ in the intermediate regime. Since the uMFT is longer in the harsh than in the mild environment (where there are fewer migration events; see Fig. 2(b)), a similar reasoning leads to a minimum mean fixation time for $\nu$ in the intermediate regime; see Fig. 4(b,d).
IV.2 Strong bottlenecks: $\psi(m,K_{-})<1$
@@
Figure 5: Typical single realisations of $N/\Omega$ (black), $N_{M}/\Omega$ (red), $N_{W}/\Omega$ (blue), and $K(t)$ (grey) vs time for different values of $K_{-}$ and $\nu$. (a,b): Here, $\nu=10^{-4}$ and $K_{-}=8$. In (a), $K=K_{-}$ at $t=0$ and $M$ and then $W$ quickly go extinct. In (b), $K=K_{+}$ at $t=0$ and $M$ fixes the population while $W$ goes extinct. (c,d): Here, $\nu=10^{-2}$ and $K_{-}=8$. In (c), mutants survive the first few bottlenecks but their abundance is low leading to the fixation of $W$ and removal of $M$ after four bottlenecks ($t\gtrsim 1000$). In (d), mutants survive the first bottlenecks and spread in the mild state where they recolonise and invade demes. They are eventually able to fix the population. (e,f): Here, $\nu=10$, and $K_{-}=4$ in (e) and $K_{-}=10$ in (f). $K(t)$ switches very frequently and is not shown for clarity. In (e), the deme size is $n\approx 2K_{-}=8$ and the dynamics is dominated by deme extinction leading to the rapid extinction of the metapopulation. In (e), the deme size is $n\approx 2K_{-}=20$ and there is $M/W$ competition that leads to fixation of $M$ and extinction of $W$ after a typical time $t\sim\theta^{\rm clique}(2K_{-})\gtrsim 10^{4}$ (not shown). Similar results are obtained on other regular graphs $\rm G$}; see text. Other parameters are $\Omega=10$, $s=0.1$, $m=10^{-4}$, and $K_{+}=200$. In all panels, initially there is a single $M$ deme and $\Omega-1$ demes occupied by $W$.
@@
Figure 6: Near-optimal condition for the idealised treatment strategy. $(\nu,K_{-})$ heatmaps of $\Phi$, $\Theta$, $\Theta_{E}$ and $(1-\Phi)(1-\Theta/\Theta_{E})$ for a clique metapopulation. The migration rate is $m=10^{-3}$ in (a-d) and $m=10^{-4}$ in (e-h). White space in panels (c) and (g) indicates where at least one realisation for those parameters did not reach extinction by $t=10^{5}$, i.e. $\Theta_{E}\sim 10^{5}$ or larger. Grey lines in panels (d) and (h) show the near-optimal conditions for the idealised treatment strategy: $\psi(m,K_{-})<1$ below the top horizontal line, $mK_{+}\theta_{E}>1$ above the bottom horizontal line, and $\nu\theta_{E}>1$ above the curved line, while the vertical line indicates where $\nu<1$ and $\theta_{E}$ from Eq. (15). The near-optimal treatment conditions is the yellowish cloud at the centre of the area enclosed by these lines. Similar results are obtained on other regular graphs $\rm G$}; see text and Fig. 10. Other parameters $\Omega=16$, $s=0.1$, and $K_{+}=200$. In all panels, initially there is a single $M$ deme and $\Omega-1$ demes occupied by $W$.
@@
When $\psi(m,K_{-})<1$, each deme can undergo strong bottlenecks; see Fig. 3(a) and below. In the harsh environment $\xi=-1$ ($K=K_{-}$), the entire metapopulation experiences extinction, in an observable time $\theta_{E}(K_{-},\Omega)$, denoted $\theta_{E}$ here for conciseness. However, in the mild state $\xi=1$ ($K=K_{+}$), deme extinction can be neglected and each site can be regarded as being occupied by either $W$ or $M$ cells. The dynamics in the harsh state is thus dominated by extinction, whereas the $M/W$ competition characterising the mild state is aptly captured by the two-state coarse-grained description of Section III.1. Environmental switching thus couples regimes that are dominated in turn by deme extinction and $M/W$ competition, yielding complex dynamical scenarios whose analysis is difficult. However, we can gain valuable insight by considering first the limits $\nu\ll 1$, $\nu\gg 1$, and then the case of intermediate switching where $\nu\lesssim 1$.
@@
When the environment varies very slowly, $\nu\ll 1$, the $K(t)$ remains at its initial value for long periods, that is $K=K_{\pm}$ if $\xi(0)=\pm 1$ each with a probability $1/2$. On the one hand, if initially $\xi=-1$ (harsh environment), $K=K_{-}$ and each deme is prone to extinction after a mean time $\tau_{E}(K_{-})$, which eventually leads to the collapse of the metapopulation after a mMET $\theta_{E}\approx e^{K_{-}}\ln{(\Omega)}/K_{-}$ when $\Omega\gg 1$ and $K_{-}\gg 1$; see Eqs. (7) and (15) and Fig. 5(a). On the other hand, if initially $\xi=1$ (mild condition), $n\approx K_{+}$ and there is $M/W$ competition characterised by the fixation of $M$ with a probability $\phi(K_{+})$ approaching 1 when $K_{+}s\gg 1$; see Eq. (14) and Fig. 5(b). As a result, when $\nu\ll 1$ and $K_{+}s\gg 1$, there are two equally likely outcomes illustrated in Fig. 5: either the extinction of the metapopulation in a mean time $\theta_{E}$ as in Fig. 5(a), or the fixation of $M$ after a mean time $\theta^{\rm G}(K_{+})$ as in Fig. 5(b).
@@
In frequently varying environments, when $\nu\gg 1$, the size of each deme readily settles about the effective carrying capacity Eq. (17), with $n\approx{\cal K}$ (when ${\cal K}\gg 1$) after $t\sim 1$ [16, 17]. Since ${\cal K}\approx 2K_{-}$ when $K_{+}\gg K_{-}$, if $\psi(m,2K_{-})<1$ the dynamics is characterised by the extinction of individual demes and then of the entire metapopulation after mean times $\tau_{E}(2K_{-})$ and $\theta_{E}(\Omega,2K_{-})$; see Fig. 5(e). However, if $\psi(m,2K_{-})\gg 1$ and $2K_{-}s\gg 1$, the dynamics is characterised by $M/W$ competition with $M$ most likely to fix the metapopulation after a mean time $\theta^{\rm G}(2K_{-})$, as illustrated by Fig. 5(f).
@@
In the intermediate switching regime, $\nu\lesssim 1$ with $\psi(m,K_{-})<1$, each deme is subject to strong bottlenecks [16, 17, 18]; see Appendix A (and first paragraph of Sec. IV). In this regime the entire metapopulation therefore experiences strong bottlenecks and can avoid extinction for extended periods of time while either strain can prevail. In the harsh environmental state ($K=K_{-}$), the dynamics is always dominated by deme extinction. In the mild state ($K=K_{+}\gg K_{-}$), there is recolonisation of empty demes that rapidly become either $W$ or $M$ demes, followed by invasions and $M/W$ competition. In order to ensure that the collapse of the metapopulation is unlikely to be observed, the mean time spent in either environmental state needs to be shorter than the metapopulation mean extinction time in the harsh environment, i.e. $1/\nu<\theta_{E}$. Moreover, when $\nu\sim 1/\tau_{E}(K_{-})>1/\theta_{E}$, numerous demes go extinct in the harsh environment before switching to the mild state. Hence, when $\nu\lesssim 1$ and $\nu\theta_{E}>1$, the metapopulation is unlikely to go extinct and transiently consists of a mixture of empty demes and $W/M$ demes before either $W$ or $M$ eventually takes over. In this regime, mutants are likely to be removed from the metapopulation when there is a small fraction of them in the harsh environment; see Fig. 5(c). At each strong bottleneck, $M$ demes have a finite probability to go extinct before switching to the mild environment, where surviving mutants can invade $W$ demes and recolonise empty demes. In a scenario illustrated by Fig. 5(c), there are periods of duration $\sim 1/\nu$ during which the number of mutants remains low and prone to extinction when $K=K_{-}$, followed by periods in $K_{+}$ where the number of $M$ demes increases (due to $M/W$ competition). Each bottleneck can thus be regarded as an attempt to remove $M$ demes, whereas each switch $K_{-}\to K_{+}$ can be envisioned as a rescue of mutants. This cycle repeats itself until $M$ demes are entirely removed after enough bottlenecks. The metapopulation thus consists of a fluctuating number of $W$ demes and empty demes. This scenario is the most likely to occur when the initial fraction of $M$ demes is small. Another possible outcome, illustrated by Fig. 5(d), occurs when mutants surviving the harsh conditions invade and are successful in recolonising many demes in the mild environment. Mutants can thus significantly increase the number of $M$ demes, exceeding that of $W$ demes. In this case, bottlenecks can be seen as an attempt to remove $W$ and $M$ demes, and the most likely outcome is the removal of $W$ demes. In this scenario, the metapopulation eventually consists of a fluctuating number of empty demes and mutant demes, as illustrated by Fig. 5(d). The results of Fig. 5 have been obtained for cliques, but the same qualitative behaviour is expected for any regular graphs $\rm G$}, with the spatial structure affecting the the long-term fraction of occupied demes and therefore the probability of removal of each bottleneck. However, phenomena operated by extinction are mostly independent of $\rm G$} and $m$, as illustrated by Fig. 10.
@@
A hypothetical idealised treatment strategy: In this intermediate switching regime, the metapopulation is likely to avoid extinction in the harsh environment if $\nu\theta_{E}\gtrsim 1$. Moreover, when $mK_{+}\theta_{E}\gtrsim 1$, then enough demes are recolonised in the mild environment to ensure that the metapopulation will not readily go extinct after a bottleneck. Hence, the metapopulation is likely to avoid extinction when $\nu\theta_{E}>1$ and $mK_{+}\theta_{E}>1$. In this scenario, either $W$ or $M$ can be entirely removed, with respective probabilities $\Phi$ and $1-\Phi$, after a mean time $\Theta$, while metapopulation extinction occurs after a mean time $\Theta_{E}\gg\Theta$. As an application, we consider a hypothetical idealised treatment to efficiently remove unwanted mutants by controlling the environmental conditions via the parameters $K_{-}$ and $\nu$. In this context, $M$ cells are interpreted as unwanted mutants that have a selective advantage over the $W$ cells we would like the population to consist of, here represented by the metapopulation consisting initially of $\Omega-1$ demes of type $W$ and a single $M$ deme. In a healthy host, cells replicate in a controlled, self-regulating manner. Mutations may lead to the loss of self-regulation of cell replication, and such mutant cells replicate rapidly. This is the case of cancerous cells, that appear as rare mutants before possibly proliferating. Thus, initial conditions like those considered here are relevant, even for smaller systems. If allowed to proliferate, cancer cells will outcompete the slower-growing healthy cells, leading to a growing tumour. Therefore, in this motivating context, we wish to remove these aggressive cells while not eliminating healthy ones. The idealised treatment strategy consists of finding the set of near-optimal environmental conditions to remove $M$ cells and minimise the risk of extinction of the entire metapopulation. This corresponds to determining the range of $K_{-}$ and $\nu$ for which $\Phi$ and $\Theta/\Theta_{E}$ are minimal. According to the above discussion, the near-optimal conditions for this idealised treatment strategy on a regular graph $\rm G$} are